ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 06 Goldstern, Martin, Technische Universitaet Wien, Austria Interpolation of monotone functions in lattices Let L be an infinite lattice. Every lattice-theoretic polynomial p(x) induces a monotone function from L to L. In which lattices is it true that every monotone function from L to L is represented or at least “approximated” (in a suitable topology) by a polynomial? I will try to give some background and discuss old and recent results.